OpenStudy (anonymous):

how to find y, if y''+2y'+y=x(e^(-x)-cosx) ?

OpenStudy (anonymous):

Have any idea? :)

OpenStudy (amistre64):

relate it to r^2 +2r +1 = 0 ; for the homogenous part

OpenStudy (amistre64):

then relate that back to some structure for the particular part

OpenStudy (amistre64):

since the characteristic equation is a perfect square; we get something resulting in: A e^rx + Bx e^rx i think

OpenStudy (anonymous):

and then ? :)

OpenStudy (mathmate):

Solve for the homogeneous case, then find the particular solution, by undetermined coefficients.

OpenStudy (anonymous):

hmm.. how to do it..?

OpenStudy (amistre64):

just since I need the practice :) our roots are r = -1,-1 \[y_h=c_1e^{-x}+c_2xe^{-x}\] if i recall it correctly, we can take this and work it for the yp \[y_p=Ae^{-x}+Bxe^{-x}\] \[y'_p=A'e^{-x}-Ae^{-x}+B'xe^{-x}-Bxe^{-x}\] we take the A and B here and equate it to 0 and down the rest \[y'_p=-Ae^{-x}-Bxe^{-x}\] \[y''_p=-A'e^{-x}+Ae^{-x}-B'xe^{-x}+Bxe^{-x}\] \[+2y'_p=-2Ae^{-x}-2Bxe^{-x}\] \[+y=Ae^{-x}+Bxe^{-x}\] ---------------------------- \[-A'e^{-x}-B'xe^{-x}=xe^{-x}-xcos(x) \] now we should reintroduce the A'B' stuff the we zeroed out and see what we can come up with

OpenStudy (amistre64):

\begin{array}l A'.e^{-x}+B'.xe^{-x}=0\\ -A'.e^{-x}-B'.xe^{-x}=x(e^{-x}-cos(x))\\ \end{array} using a cramer rule we can see what happens. \[B'=\frac{xe^{-x}(e^{-x}-cos(x))}{-xe^{-2x}+xe^{-2x}}\] ack!!, looks like i ended up with a zero in a denominator.

OpenStudy (amistre64):

undetermined coeffs fer sure

OpenStudy (amistre64):

this might be more useful than the stuff im trying to remember :) http://www.wolframalpha.com/input/?i=+y%27%27%2B2y%27%2By%3Dx%28e%5E%28-x%29-cosx%29

OpenStudy (amistre64):

ahh, it took and split up the particulate solution into xe^-x and -xcos(x) so we had to work it out twice

OpenStudy (anonymous):

good work